40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue May 20 02:12:35 2014, Jenne, Update, LSC, ALS X noise investigation ALS_outofloop_19May2013.pdfXbeatSaturate.png
    Reply  Wed May 21 13:20:34 2014, manasa, Update, LSC, ALS X noise from angular motion of mirrors ALSX_OLYerrITM.png
       Reply  Thu May 22 00:30:40 2014, manasa, Update, LSC, ALS X noise from angular motion of mirrors ALSX_OLPerrITM2.pngALSX_OLYerrITM.pngALSX_OLPerrETM.pngALSX_OLYerrETM2.png
Message ID: 9983     Entry time: Wed May 21 13:20:34 2014     In reply to: 9972     Reply to this: 9988
Author: manasa 
Type: Update 
Category: LSC 
Subject: ALS X noise from angular motion of mirrors 

Quote:

[Rana, Jenne]

9.  Arm cavity alignment.  Significant DC effect.

    * When the alignment of one of the arm cavity mirrors is changed, the DC value of the beatnote signal changes. 

           * ITMX moved in yaw, we see a 7kHz/15urad DC shift in the BEATX_FINE_PHASE_OUT_HZ time series.

          * ETMX moved in yaw, we see an 8kHz/5.5urad DC shift in the time series.  We aren't sure why this is about a factor of 3 times larger effect (same shift for smaller misalignment) than the ITM.

    * We want to do a Yuta-style analysis to see what the angle to length coupling looks like, so that we can measure the angular motion of our cavity mirrors and put the expected noise into our ALS noise budget.  Perhaps this will help us understand the low frequency difference between our in-loop beatnote error signal and our in-loop PDH error signals (red vs. maroon on the ALS noise budget posted above Pianosa). 

    * I've asked Manasa to take some transfer functions in the morning, so that we can start to have an idea of what is going on with this.

Attached is the measurement of the transfer function from ITMX oplev error in yaw to the ALSX error signal.

The arm was locked to the IR using POX and the green beat frequency (between X arm trans in green and PSL green) in this case was 27MHz.

The transfer function looks mostly flat between 1Hz - 30Hz at 700-800 Hz/urad. The DC shift that Jenne measured from the time series is ~500 Hz/urad.

So far I have not been able to measure the TF below 1Hz without the arm losing its lock. Updates will follow.

Data xml file can be found in /users/manasa/data/140521/

Attachment 1: ALSX_OLYerrITM.png  57 kB  | Hide | Hide all
ALSX_OLYerrITM.png
ELOG V3.1.3-